A and B are 2 vectors such that their resultant is perpendicular to A and magnitude is equal to A find the magnitude of vector B

Grade:11

1 Answers

Shashank Singh
6 years ago
(A+B).A=0 (Since they are perpendicular)
IAI^2 + B.A=0
B.A= -IAI^2
IA+BI= IAI
IAI^2+IBI^2+2A.B=IAI^2
IAI^2+IBI^2- 2IAI^2= IAI^2
IBI^2= 2IAI^2
IBI= Sqrt 2IAI